Jane has 12.00 less than twice the amount that Jason saved. Together they saved 144.00. How much did Jane save? how much did Jason save?
Answer by JulietG(1812)

N + S = 144 (Together they saved 144.00.)
N = 2S-12 (Jane has 12.00 less than twice the amount that Jason saved.)
Substitute the value of N from the second equation into the first.
(2S-12) + S =144
Add
3S-12=144
Add 12 to each side
3S = 156
Divide each side by 3
S = 52
.
If Jason saved 52, then Jane saved (52*2) - 12. 104-12 = 92. 92+52=144.
Success!